# for Generalist Vision-Language-Action Policy

### v1

@@ 3.1 Architecture
Model Design. Our core design is the DiT structure [56], which denoises action token chunks rather than individual action tokens. This is achieved by conditioning directly on image observations and instruction tokens through an in-context conditioning approach using a causal Transformer. Specifically, we concatenate language tokens, image features, and timestamp embeddings at the beginning of the sequence, treating the noisy action in conjunction with the instruction tokens, as illustrated in Figure 2. This design preserves the scalability of Transformer networks and enables denoising to be conditioned directly on image patches, thereby allowing the model to capture nuanced changes in action over historical observations. The model is supervised by the noise introduced into the continuous actions. In other words, we directly apply the diffusion objective in the action chunk space with a large Transformer model, in contrast to the diffusion action head approach [72, 18, 61]. Notably, our proposed Dita presents a versatile and scalable design, adaptable to diverse datasets for both pretraining and finetuning, while achieving promising performance. Furthermore, additional observation tokens and input can be seamlessly integrated into the Transformer architecture. Further details are provided in Appendix A.
@@ 3.2 Training Objective
The denoising network $\mathcal{E}_{\theta}(c_{lang},c_{obs},t,{\bm{x}}^{t})$ is constructed upon a causal Transformer, where $c_{obs}$ represents the image observation, $c_{lang}$ denotes the language instruction, and $t\in{1,2,\dots,T_{train}}$ is the timestamp index within the total denoising steps $T_{train}$. During training, a Gaussian noise vector ${\bm{x}}^{t}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$ is sampled at each timestamp $t$ and added to the action ${\bm{a}}$ to form the noised action token $\hat{{\bm{a}}}$. The network $\mathcal{E}_{\theta}$ is trained to predict the noise vector $\hat{{\bm{x}}}$, with randomly sampled $t$ The optimization objective of Dita is to minimize the mean squared error (MSE) loss between ${\bm{x}}^{t}$ and $\hat{{\bm{x}}}^{t}$.
@@ 3.4 Pretraining Details
We employ the DDPM diffusion objective [25] with $T_{train}=1000$ timestamps for pretraining, while adopting DDIM [69] with $T_{eval}=20$ timestamps during zero-shot evaluation to accelerate inference. Based on preliminary experiments reported in ManiSkill2 [22], we utilize 2-frame image observations to predict 16 action chunks. The network is optimized by AdamW [43] for 100,000 steps, with learning rates of $1e{-4}$ for both the causal Transformer and Q-Former, and $1e{-5}$ for DINOv2. Training is conducted with a batch size of 8192 across 32 NVIDIA A100 GPUs, allocating 256 samples per GPU. Additional pretraining configurations are detailed in Appendix A.
@@ 4 Simulation Experiments
Method coke__\__can move__\__near drawer
@@ 4.5 ManiSkill2
Setup. To construct the benchmark, we select 5 tasks (PickCube-v0, StackCube-v0, PickSingleYCB-v0, PickClutterYCB-v0, PickSingleEGAD-v0) from ManiSkill2 and create a camera pool comprising 300K random cameras. 20 cameras are sampled each time to render each trajectory, resulting in over 40K trajectories, which are utilized to train Dita from scratch. The generated dataset is divided into training and validation sets with a 19:1 ratio, ensuring that each category in task family, and trajectories rendered from different camera views are assigned to both, thereby preventing data leakage. During training, the number of data samples is balanced across task families by duplicating trajectories for task families with fewer samples. To construct a closed-loop evaluation dataset, we randomly sample 100 trajectories from the validation set for each task family. This evaluation dataset with 500 trajectories is used to assess the success rate for each task family and demonstrate the camera-view generalization capabilities of Dita. We also implement RT-1 [8] style baseline model $\mathcal{E}_{\theta\sim s}^{Disc}$ with an architecture similar to ours for comparison. Unlikely, we discretize each action dimension into 256 bins [8] and utilize a Transformer network to predict the corresponding bin indices.
@@ 4.6 Ablation Study
# Steps 100 50 20 10 5 2
Pick Coke (variant) 76.4 79.1 85.5 85.3 82.7 70.4
Pick Coke (match) 79.7 83.3 83.7 82.0 82. 73.3
Move Near (variant) 52.1 66.0 73.0 69.5 63.5 51.6
Move Near (match) 49.1 72.0 76.0 74.0 72.0 65.0
@@ 5.1 Real-Robot Task Finetuning
Setup. To deploy the Dita model, as shown in Figure 3, the robot setup consists of a 7-DoF tabletop Franka Emika Panda robot arm and a Robotiq 2F-85 gripper. A RealSense D435i RGB-D camera, positioned approximately 1.5m away from the robot in a third-person view, captures RGB scenes with cluttered background at each inference timestamp. Robot control is managed from a desktop computer running ROS, communicating with the model-deploy server with 1 NVIDIA A100 GPU. The system is operating under control frequency of 3Hz. Given the data domain gap between our robot platform and the pretrain dataset, we primarily evaluate Dita on 10-shot generalization for the following challenging tasks relevant to current VLA approaches:
@@
Optimization. We finetune the Dita on the aforementioned multiple manipulation tasks, with data collected on the same platform, using LoRA [26] for fair comparison and AdamW for 20,000 steps with image augmentations. The number of timestamps is set to 100 for DDPM [25], and the batch size of 512.
@@ 6 Conclusion
In this paper, we present Dita, an architecture for generalist robot learning that leverages a Transformer-based diffusion model to denoise continuous action sequences through an in-context conditioning mechanism. By harnessing the scalability of Transformers, Dita effectively models diverse robot behaviors across extensive cross-embodiment datasets, enabling robust generalization across multiple simulation benchmarks within a unified framework. Additionally, Dita demonstrates strong few-shot adaptation capabilities, successfully transferring to novel real-world robot setups and long-horizon tasks with minimal in-domain samples. Notably, the model is clean, lightweight, and open-source, and its promising performance—achieved exclusively with a single third-person camera input—underscores its potential as a scalable and flexible solution for learning generalist robot policies.
@@ B.2 LIBERO
Optimization. We optimize the network using AdamW for 100,000 steps on LIBERO. The learning rate is set to $41e-$ 4 for LIBERO-SPATIAL, LIBERO-OBJECT, and LIBERO-GOAL, and $45e-$ 4 for LIBERO-LONG. Across all sub-datasets, a half-cycle cosine scheduler is applied to decay the learning rate. Denoising timestamps are set to 100 during finetuning, and training is conducted with a batch size of 512 across 8 NVIDIA A100 GPUs.
@@ B.3 CALVIN
Optimization. For each training iteration, the model predicts 10 future action chunks supervised by MSE loss. An AdamW optimizer is used together with a decayed learning rate with half-cycle cosine scheduler after several steps of warming up. The learning rate is initialized as $41e-$ 4. Model is trained for 15 epochs with batch size of 128 across 4 NVIDIA A100 GPUs.
@@ B.4 Maniskill2
Optimization. The network is optimized using AdamW for 50,000 steps on ManiSkill2, with a learning rate set to $41e-$ 4. The number of denoising timestamps is set to 100, and the batch size is 1024 distributed across 16 NVIDIA A100 GPUs.
@@ D.2 Convergence Analysis
D.3 More Ablations
Learning Rate Scheduler. As outlined in the main text, we utilize a standard learning rate scheduler to decay the learning rate during the experiments in Calvin, rather than using a fixed learning rate of $41e-$ 4 as in the pretraining stage. This adjustment results in a slight performance improvement, as shown in Table 8.
@@ D.3 More Ablations
Shuffle Buffer Size coke__\__can

### v2

@@ 3.1 Architecture
Model Design. Our core design is the DiT structure [56], which denoises the action chunk with multiple action tokens. This is achieved by conditioning directly on image observations and instruction tokens through an in-context conditioning approach using a causal Transformer. Specifically, we concatenate language tokens, image features, and timestep embeddings at the beginning of the sequence, treating the noisy action in conjunction with the instruction tokens, as illustrated in Figure 2. This design preserves the scalability of Transformer networks and enables denoising to be conditioned directly on image patches, thereby allowing the model to capture nuanced changes in action over historical observations. The model is supervised by the noise introduced into the continuous actions. In other words, we directly apply the diffusion objective in the action chunk space with a large Transformer model, in contrast to the diffusion action head approach [72, 18, 61]. Notably, our proposed Dita presents a versatile and scalable design, adaptable to diverse datasets for both pretraining and finetuning, while achieving promising performance. Furthermore, additional observation tokens and input can be seamlessly integrated into the Transformer architecture. Further details are provided in Appendix A.
@@ 3.2 Training Objective
The denoising network $\mathcal{E}_{\theta}(c_{lang},c_{obs},t,{\bm{x}}^{t})$ is constructed upon a causal Transformer, where $c_{obs}$ represents the image observation, $c_{lang}$ denotes the language instruction, and $t\in{1,2,\dots,T_{train}}$ is the timestep index within the total denoising steps $T_{train}$. During training, a Gaussian noise vector ${\bm{x}}^{t}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$ is sampled at each timestep $t$ and added to the action $\bm{a}$} to form the noised action token $\hat{{\bm{a}}}$. The network $\mathcal{E}_{\theta}$ is trained to predict the noise vector $\hat{{\bm{x}}}$, with randomly sampled $t$ . The optimization objective of Dita is to minimize the mean squared error (MSE) loss between ${\bm{x}}^{t}$ and $\hat{{\bm{x}}}^{t}$.
@@ 3.4 Pretraining Details
We employ the DDPM diffusion objective [25] with $T_{train}=1000$ timesteps for pretraining, while adopting DDIM [69] with $T_{eval}=20$ timesteps during zero-shot evaluation to accelerate inference. Based on preliminary experiments reported in ManiSkill2 [22], we utilize 2-frame image observations to predict 16 action chunks. The network is optimized by AdamW [43] for 100,000 steps, with learning rates of 1e41e{-4} for both the causal Transformer and Q-Former, and 1e51e{-5} for DINOv2. Training is conducted with a batch size of 8192 across 32 NVIDIA A100 GPUs, allocating 256 samples per GPU. Additional pretraining configurations are detailed in Appendix A.
@@ 4 Simulation Experiments
Method coke_\_can move_\_near drawer
@@ 4.5 ManiSkill2
Setup. To construct the benchmark, we select 5 tasks (PickCube-v0, StackCube-v0, PickSingleYCB-v0, PickClutterYCB-v0, PickSingleEGAD-v0) from ManiSkill2 and create a camera pool comprising 300K random cameras. More details are provided in Appendix. We also implement RT-1 [8] style baseline model $\mathcal{E}_{\theta\sim s}^{Disc}$ with an architecture similar to ours for comparison. Unlikely, we discretize each action dimension into 256 bins [8] and utilize a Transformer network to predict the corresponding bin indices.
@@ 5.1 Real-Robot Task Finetuning
Setup. To deploy the Dita model, as shown in Figure 3, the robot setup consists of a 7-DoF tabletop Franka Emika Panda robot arm and a Robotiq 2F-85 gripper. A RealSense D435i RGB-D camera, positioned approximately 1.5m away from the robot in a third-person view, captures RGB scenes with cluttered background at each inference timestep. Robot control is managed from a desktop computer running ROS, communicating with the model-deploy server with 1 NVIDIA A100 GPU. The system is operating under control frequency of 3Hz. Given the data domain gap between our robot platform and the pretrain dataset, we primarily evaluate Dita on 10-shot generalization for the following challenging tasks relevant to current VLA approaches:
@@
Optimization. We finetune the Dita on the aforementioned multiple manipulation tasks, with data collected on the same platform, using LoRA [26] for fair comparison and AdamW for 20,000 steps with image augmentations. The number of timesteps is set to 100 for DDPM [25], and the batch size of 512.
@@ 6 Conclusion
We present Dita, an architecture for generalist robot learning that leverages a Transformer-based diffusion model to denoise continuous action sequences through an in-context conditioning mechanism. By harnessing the scalability of Transformers, Dita effectively models diverse robot behaviors across extensive cross-embodiment datasets, enabling robust generalization across multiple simulation benchmarks within a unified framework. Additionally, Dita demonstrates strong few-shot adaptation capabilities, successfully transferring to novel real-world robot setups and long-horizon tasks with minimal in-domain samples. Notably, the model is clean, lightweight, and open-source, and its promising performance—achieved exclusively with a single third-person camera input—underscores its potential as a scalable and flexible solution for generalist policies.
7 Acknowledgment
This work is supported by the National Key R&D Program of China (NO. 2022ZD0160201).
@@ B.2 LIBERO
Optimization. We optimize the network using AdamW for 100,000 steps on LIBERO. The learning rate is set to 1e41e-4 for LIBERO-SPATIAL, LIBERO-OBJECT, and LIBERO-GOAL, and 5e45e-4 for LIBERO-LONG. Across all sub-datasets, a half-cycle cosine scheduler is applied to decay the learning rate. Denoising timestamps are set to 100 during finetuning, and training is conducted with a batch size of 512 across 8 NVIDIA A100 GPUs.
@@ B.3 CALVIN
Optimization. For each training iteration, the model predicts 10 future action chunks supervised by MSE loss. An AdamW optimizer is used together with a decayed learning rate with half-cycle cosine scheduler after several steps of warming up. The learning rate is initialized as 1e41e-4. Model is trained for 15 epochs with batch size of 128 across 4 NVIDIA A100 GPUs.
@@ B.4 Maniskill2
Setup. To construct the benchmark, we select 5 tasks (PickCube-v0, StackCube-v0, PickSingleYCB-v0,PickClutterYCB-v0, PickSingleEGAD-v0) from ManiSkill2 and create a camera pool comprising 300K random cameras. 20 cameras are sampled each time to render each trajectory, resulting in over 40K trajectories, which are utilized to train Dita from scratch. The generated dataset is divided into training and validation sets with a 19:1 ratio, ensuring that each category in task family, and trajectories rendered from different camera views are assigned to both, thereby preventing data leakage. During training, the number of data samples is balanced across task families by duplicating trajectories for task families with fewer samples. To construct a closed-loop evaluation dataset, we randomly sample 100 trajectories from the validation set for each task family. This evaluation dataset with 500 trajectories is used to assess the success rate for each task family and demonstrate the camera-view generalization capabilities of Dita.
Optimization. The network is optimized using AdamW for 50,000 steps on ManiSkill2, with a learning rate set to 1e41e-4. The number of denoising timestamps is set to 100, and the batch size is 1024 distributed across 16 NVIDIA A100 GPUs.
@@ D.1 Practices on reproducing Octo and OpenVLA
[…]
Method No. Instructions in a Row (1000 chains)
1 2 3 4 5 Avg.Len.
Dita w/o aug 94.5% 82.5% 72.8% 61.3% 50.0% 3.61
Dita w/aug 92.3% 83.8% 75.8% 67.3% 69.0% 3.78
w/o pretraining w/ aug
Dita (2 layers) 75.8% 47.3% 29.3% 19.0% 12.5% 1.84
Dita (6 layers) 80.5% 60.0% 39.0% 26.3% 15.0% 2.21
Dita (12 layers) 84.5% 65.5% 48.8% 34.5% 22.5% 2.56
Dita (24 layers) 87.8% 67.5% 48.5 % 34.5 % 23.0% 2.61
Dita 84.5% 65.5% 48.8% 34.5% 22.5% 2.56
AdaLN DiT [17] 68.3% 40.0% 21.0% 11.3% 5% 1.45
@@ D.2 Convergence Analysis
D.3 Failure Analysis
Figure 11 illustrates two representative failure cases. The first case involves a failed grasp during execution. Although the model is capable of recovering and retrying, the failure suggests an inability to approach the correct position for a successful pickup. The second case concerns pouring while in motion, where the model tends to spill the contents due to unstable handling.
Figure 11: Failure Cases Analysis.
@@ D.3 Failure Analysis
D.4 More Ablations
Learning Rate Scheduler. As outlined in the main text, we utilize a standard learning rate scheduler to decay the learning rate during the experiments in Calvin, rather than using a fixed learning rate of 1e41e-4 as in the pretraining stage. This adjustment results in a slight performance improvement, as shown in Table 8.
@@ D.4 More Ablations
Shuffle Buffer Size coke_\_can
@@
# Steps 100 50 20 10 5 2
Pick Coke (variant) 76.4 79.1 85.5 85.3 82.7 70.4
Pick Coke (match) 79.7 83.3 83.7 82.0 82. 73.3
Move Near (variant) 52.1 66.0 73.0 69.5 63.5 51.6
Move Near (match) 49.1 72.0 76.0 74.0 72.0 65.0